Suppose the information in the following table is for a simple economy that produces only four goods and services: shoes, hamburgers, shirts, and cotton. Assume that shoes, hamburgers, and shirts are final goods and that all the cotton is used in the production of shirts. Calculate the GDP deflator for 2015. Use 2009 as the base year for the calculations. Provide your answer as a number rounded to two decimal places. Do not include any symbols, such as "$," "=," "%," or "," in your answer. 2009 2014 2015 Product Quantity Price Quantity Price Quantity Price Shoes 75 $45 95 $59 114 $67 Hamburgers 73 2 109 4 96 4 Shirts 50 30 70 43 70 38 Cotton 50 7 70 12 70 18
Answer #1

Since cotton is the intermediate good it will not be counted in GDP

Hence Nominal GDP of year 2015 = Shoes Price 2015* Shoes Quantity 2015 +Hamburger Price 2015* Hamburger Quantity 2015+Shirts Price 2015* Shirts Quantity 2015

= 67*114+4*96+38*70 = 7638+384+2660=10682

Real GDP of year 2015 = Shoes Price 2009* Shoes Quantity 2015 +Hamburger Price 2009* Hamburger Quantity 2015+Shirts Price 2009* Shirts Quantity 2015

= 45*114+2*96+30*70 = 5130+192+2100=7422

GDP Deflator = Nominal GDP/ Real GDP *100 = 10682/7422 *100 = 1.4392*100=143.92